OpenStudy (anonymous):

IIIIIMM BAAAAAACK!!!!!!!!! [5.02] Given parallelogram LMNO with the diagonals intersecting at point P. If MP = 2x - 1 and OP = 3x - 4, find OP. (4 points) -1 1 3 5

OpenStudy (igreen):

OP = 3x - 4 Find OP ?

OpenStudy (anonymous):

1 im pretty sure

OpenStudy (anonymous):

u posotive it one?

OpenStudy (anonymous):

YEA

OpenStudy (anonymous):

sweet thx

OpenStudy (igreen):

No 1 is incorrect.

OpenStudy (igreen):

First we find 'x': 2x - 1 = 3x - 4 Add 1 to both sides: 2x = 3x - 3 Subtract 3x to both sides: -1x = -3 Divide -1 to both sides, what do you get?

OpenStudy (anonymous):

wat is it then 3?

OpenStudy (anonymous):

ohh i miss read sorry i dont have my glasses on

OpenStudy (anonymous):

i hate that!

OpenStudy (igreen):

Yes, we get 3..but that's not our answer..now we have to plug in 3 into the equation for OP. 3x - 4 3(3) - 4 Solve that and you'll get your answer.

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

it would be 5 right cuz 3*3=9-4=5

OpenStudy (igreen):

No.. @countryboy212 that's wrong and please don't give answers.

OpenStudy (igreen):

Yep, you got it. @WhiteFangIII

OpenStudy (anonymous):

ay